Assume that a student did not have time to study and is simply guessing at the corect answer foc question (a) What percentage of the questions would you expect the student to answer coety (b) What is the standard deviation of the percentage of questions the student will answer corectny
Answer #1

here this is binomial distribution with parameter n=25 and p=probability of getting a answer correct =1/5=0.2

a) hence % of question answered correctly =p*100=(0.2)*100=20%

b)std deviation =sqrt(p*(1-p)/n)*100=sqrt(0.2*0.8/25)*100=8.00%